OpenStudy (anonymous):

I'd be really happy someone would help me. Explain how to add unlike fractions. It says to use the largest denominator as the common denominator. 3/4+1/8

OpenStudy (skullpatrol):

Any ideas @anaerrez , since you need ideas to get answers...

OpenStudy (anonymous):

Unfortunately no, I've been struggling for day's and i was hoping someone culd explain this as simple as possible.

OpenStudy (skullpatrol):

have you covered least common denominators?

OpenStudy (amistre64):

spose we have 2 candy bars, one is broken into 4 pieces, the other into 8 pieces if we want to evaluate each candy bar on the same basis, they both need to be broken into the same number of parts ... finding a common denominator

OpenStudy (skullpatrol):

@anaerrez are you still there?

OpenStudy (amistre64):

can we break the candy bars such that they are each broken into the same nuimber of pieces? can 4 parts be broken up nicely to make 8 even parts? can 8 parts be broken up nicely to make 4 even parts? if not, then can we devise a way to break them evenly until they are in an equal number of parts?

OpenStudy (anonymous):

Yes

OpenStudy (skullpatrol):

The first rule you need to use is the multiplication rule for fractions: $$\huge\frac{c}{d}=\frac{bc}{bd}$$

OpenStudy (anonymous):

Ok i dont understand. I get that im supposed to rewrite the problem using the largest denominator as the common denominator but where do I go from there.?

OpenStudy (skullpatrol):

$$\huge\frac {3}{4}+\frac{1}{8}=\huge\frac {3*2}{4*2}+\frac{1}{8}$$

OpenStudy (skullpatrol):

$$\huge\frac {6}{8}+\frac{1}{8}=\huge\frac {6+1}{8}$$